OpenStudy (anonymous):

Amelia rolls two fair number cubes numbered from 1 to 6. She first defines the sample space as shown below: (1, 1), (1, 2), (1, 3), (1, 4), (1, 5), (1, 6) (2, 1), (2, 2), (2, 3), (2, 4), (2, 5), (2, 6) (3, 1), (3, 2), (3, 3), (3, 4), (3, 5), (3, 6) (4, 1), (4, 2), (4, 3), (4, 4), (4, 5), (4, 6) (5, 1), (5, 2), (5, 3), (5, 4), (5, 5), (5, 6) (6, 1), (6, 2), (6, 3), (6, 4), (6, 5), (6, 6) Based on the sample space, what is the probability of getting a total of 4? 6 over 36 5 over 36 4 over 36 3 over 36

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

someone c:

OpenStudy (lynfran):

4/36

OpenStudy (lynfran):

o wait let me recheck..

OpenStudy (anonymous):

ooooo mkay

OpenStudy (lynfran):

i dont really understand this part Based on the sample space, what is the probability of getting a total of 4? getting 4 #??

OpenStudy (lynfran):

o i got it...

OpenStudy (lynfran):

ok so you need to look at the sample space and add the 2value in the brackets...they must add up to 4...so far i see thiss.... (1,3) 1+3=4 (2,2) 2+2=4 (3,1) 3+1=4 so the probability is 3/36